# On Nanocones as a Gravitational Analog System

### v1

@@ VI Final considerations
In this article, we delve into the analogies between material physics and the structure of spacetime. Several analogies between condensed matter physics and gravitation can be found in the literature kraus2019unexpected ; fagnocchi2008analog . We were able to construct a topological defect in spacetime and an equivalent one on graphene and BN monolayers. From the energy definition of TEGR, we predicted a linear response to the disclination angle in the total energy of spacetime. By identifying this total energy with the geometrical (torsional) energy of the nanocones, we found that the definitions of TEGR can be used to evaluate the torsional energy of the material if the constant, analogous to the gravitational coupling one, of the material is determined. We conjecture that these results are not coincidental but part of a greater result, i.e., the energy expression (39) is valid not only for sheets of any material but for all material structures. This conjecture, together with the known analogies between matter and gravity, is a strong indication that the description of gravity by TEGR is more precise than GR. Since the dynamics of TEGR are equivalent to those of GR, all dynamical analogies between GR and material science are valid in TEGR. However, the absence of a consistent definition of gravitational energy in GR makes the results presented here impossible to obtain in Einstein’s theory. Therefore, the results of this article strongly favor TEGR. Given that graphene nanocones are not merely theoretical structures, since they have already been synthesized experimental , further experimental investigations could verify the present theoretical results.
@@
The main results presented in Figures 3 and 4 show small deviations from a perfectly linear behavior independent of the size of the structure. It is worth noting that the Universal Force Field theory presented in section IV is a good approximation compared to experimental results, i.e., it contains several classical approximations to a quantum problem. Without these approximations the evaluation would be virtually impossible for huge molecular systems like the nanocones. Besides, it is a numerical solution, and small deviations are to be expected. Hence, similar to experimental data, simulation data contains an uncertainty that cannot be removed. In addition to these, we have four other causes for the errors.
@@
First, the model constructed in section III considers a circular disclination core with radius $r=a$. In the limit $a\rightarrow\infty$, we have a disclination core along the $z$ axis only, i.e., a cosmic string. In the real case of the structures, the hexagonal pattern of its atoms makes it impossible to have a punctual disclination core, even a circular one with a small radius. In order to mitigate these effects, we considered structures with an intermediate size, avoiding structures with a small radius. For structures with a diameter around eight hexagons, the correlation to the linear behavior was smaller. The size of the structure does not carry out to the core; thus, with intermediate structures, the contribution of these regions is smaller in comparison. One may think that this contribution is the same for the same sheet, but in reality, the configuration of the disclination core changes with the disclination angle. Let us consider Sheet 1, for example. The core(peak) of the nanocone for $\Delta\phi=60^{\circ}$ can be seen in Figure 5.
@@
For $\Delta\phi=180^{\circ}$ and $\Delta\phi=240^{\circ}$, as shown in Figures 7 and 8 respectively, we observe three and four pentagons forming the core.
@@
Finally, the Universal Force Field (UFF) is classical, meaning they is non-relativistic. As a result, the UFF cannot accurately represent materials under relativistic dynamics without introducing errors. This limitation is inherent in its design, as it does not account for relativistic effects.
Given the significance of the qualitative results, we are prompted to ask: how can we quantitatively validate expression (39)? This validation hinges on the nature of the coupling constant $k$. If we interpret spacetime as a material, this constant is associated with some of its internal properties. In SI units, $ck$ has the dimension of force. To associate it with the sheet, we must find an internal characteristic of the material represented by $ck$. The Young modulus of the material could partially express this constant, but it yields a magnitude order distinct from the simulation results. Another possibility is the interatomic force between the atoms in the structure, which already has the right dimension. Finding this force is challenging, but we can evaluate the bound energy between the atoms, i.e., the atom-atom electronic interactions. With the bound energy, we can estimate the force by dividing it by the interatomic mean distance. This yields a rough estimate for the force, albeit with the correct magnitude order.
In order to conduct the simulation, we employed another molecular theory, Density Functional Theory (DFT). DFT is an ab initio theory that allows for the study of molecular dynamics by describing the electrons as an electronic density field. Unlike the Universal Force Field (UFF), it is a quantum approach to the problem. Since the theoretical background of this theory is extensive and these simulations are not the main results of our article, we refer the reader to the literature for detailed information parr1983density ; gross2013density .
@@
$\alpha_{C}=2kL=10.452\,eV\,,$ (59)
where $L=140\,pm$ (approximate diameter of a carbon atom) is the thickness of the sheet. We can compare (59) with the results of the linear fit (45-48). This rough calculation provides a very good approximation with the correct magnitude order and even a close numerical match for Sheet 4, where we obtained $\alpha=10.516\,eV$. Notably, larger sheets yielded more accurate results.
@@
$\alpha_{BN}=2kL=9.965\,eV\,.$ (62)
Again, we observe the same magnitude order and a very good numerical approximation for Sheet BN 6, where $\alpha=9.226\,eV$. Additionally, we expect lower general values for the coefficient in the BN sheet, as the force between atoms is slightly smaller, reflected in the smaller values in equations (49-54).
If the preceding hypothesis is accurate, as substantiated by these findings, and we conceptualize spacetime as a cellular structure comprised of minuscule particles, we can deduce that the force between these particles corresponds to the coupling constant. Expressed in SI units, this equation results in $c\,k=\frac{c^{4}}{16\pi G}=2.41596\times 10^{42}N$. The immense magnitude of this force indicates that spacetime possesses an extremely high resistance to deformation (about $10^{30}$ times greater than that of a one-square-meter graphene sheet). Consequently, gravity, which arises from the deformation of spacetime, is a remarkably feeble phenomena if the force of deformation is insufficient, meaning that smaller amounts of energy cannot cause considerable deformation of spacetime. This aligns with our understanding of gravity.
Acknowledgements
The authors are grateful to Professor J. W. Maluf for his invaluable assistance in revising the manuscript and for providing insightful ideas that greatly contributed to this work.

### v2

@@ VI Final considerations
In this article, we delve into the analogies between material physics and the structure of spacetime. Several analogies between condensed matter physics and gravitation can be found in the literature kraus2019unexpected ; fagnocchi2008analog . We were able to construct a topological defect in spacetime and an equivalent one on graphene and BN monolayers. From the energy definition of TEGR, we predicted a linear response to the disclination angle in the total energy of spacetime. By identifying this total energy with the geometrical (torsional) energy of the nanocones, we found that the definitions of TEGR can be used to evaluate the torsional energy of the material if the constant, analogous to the gravitational coupling one, of the material is determined. We conjecture that these results are not coincidental but part of a greater result, i.e., the energy expression (39) is valid not only for sheets of any material but for all material structures. This conjecture, together with the known analogies between condensed matter physics and gravity, is a strong indication that the description of gravity by TEGR is more precise than GR. Since the dynamics of TEGR are equivalent to those of GR, all dynamical analogies between GR and material science are valid in TEGR. However, the absence of a consistent definition of gravitational energy in GR makes the results presented here impossible to obtain in Einstein’s theory. Therefore, the results of this article strongly favor TEGR. Given that graphene nanocones are not merely theoretical structures, since they have already been synthesized experimental , further experimental investigations could verify the present theoretical results.
@@
The main results presented in Figures 4 and 5 show small deviations from a perfectly linear behavior independent of the size of the structure. It is worth noting that the UFF theory presented in section IV is a good approximation compared to experimental results, i.e., it contains several classical approximations to a quantum problem. Without these approximations the evaluation would be virtually impossible for huge molecular systems like the nanocones. Besides, it is a numerical solution, and small deviations are to be expected. Hence, similar to experimental data, simulation data contains an uncertainty that cannot be removed. Moreover, we have four other causes for the errors.
@@
First, the model constructed in section III considers a circular disclination core with radius $r=a$. In the limit $a\rightarrow 0$, we have a disclination core along the $z$ axis only, i.e., a cosmic string. In the real case of the structures, the hexagonal pattern of its atoms makes it impossible to have a punctual disclination core, even a circular one with a small radius. In order to mitigate these effects, we considered structures with an intermediate size, avoiding structures with a small radius. For structures with a diameter around eight hexagons, the correlation to the linear behavior was smaller. The size of the structure does not carry out to the core; thus, with intermediate structures, the contribution of these regions is smaller in comparison. One may think that this contribution is the same for the same sheet, but in reality, the configuration of the disclination core changes with the disclination angle. Let us consider Sheet 1, for example. The core(peak) of the nanocone for $\Delta\phi=60^{\circ}$ can be seen in Figure 6.
@@
For $\Delta\phi=180^{\circ}$ and $\Delta\phi=240^{\circ}$, as shown in Figures 8 and 9 respectively, we observe three and four pentagons forming the core.
@@
Finally, the UFF is non-relativistic. As a result, the UFF cannot accurately represent materials under relativistic dynamics without introducing errors. This limitation is inherent in its design, as it does not account for relativistic effects.
Given the significance of the qualitative results, we are prompted to ask: how can we quantitatively validate expression (39)? This validation hinges on the nature of the coupling constant $\kappa$. If we interpret spacetime as a material, this constant is associated with some of its internal properties. In SI units, $c\,\kappa$ has the dimension of force. To associate it with the sheet, we must find an internal characteristic of the material represented by $c\,\kappa$. The Young modulus of the material could partially express this constant, but it yields a magnitude order distinct from the simulation results. Another possibility is the interatomic force between the atoms in the structure, which already has the right dimension. Finding this force is challenging, but we can evaluate the bound energy between the atoms, i.e., the atom-atom electronic interactions. With the bound energy, we can estimate the force by dividing it by the interatomic mean distance. This yields a rough estimate for the force, albeit with the correct magnitude order.
In order to conduct the simulation, we employed another molecular theory, Density Functional Theory (DFT). DFT is an ab initio theory that allows for the study of molecular dynamics by describing the electrons as an electronic density field. Unlike the UFF, it is a quantum approach to the problem. Since the theoretical background of this theory is extensive and these simulations are not the main results of our article, we refer the reader to the literature for detailed information parr1983density ; gross2013density .
@@
$\alpha_{C}=2c\kappa L=10.452\,eV\,,$ (57)
where $L=140\,pm$ (approximate diameter of a carbon atom) is the thickness of the sheet. We can compare (57) with the results of the linear fit (43-46). This rough calculation provides a very good approximation with the correct magnitude order and even a close numerical match for Sheet 4, where we obtained $\alpha=10.516\,eV$. Notably, larger sheets yielded more accurate results.
@@
$\alpha_{BN}=2c\kappa L=9.965\,eV\,.$ (60)
Again, we observe the same magnitude order and a very good numerical approximation for Sheet BN 6, where $\alpha=9.226\,eV$. Additionally, we expect lower general values for the coefficient in the BN sheet, as the force between atoms is slightly smaller, reflected in the smaller values in equations (47-52).
If the preceding hypothesis is accurate, as substantiated by these findings, and we conceptualize spacetime as a cellular structure comprised of minuscule particles, we can deduce that the force between these particles corresponds to the coupling constant. Expressed in SI units, this equation results in $c\,\kappa=\frac{c^{4}}{16\pi G}=2.41596\times 10^{42}N$. The immense magnitude of this force indicates that spacetime possesses an extremely high resistance to deformation (about $10^{30}$ times greater than that of a one-square-meter graphene sheet). Consequently, gravity, which arises from the deformation of spacetime, is a remarkably feeble phenomena. If the force of deformation is insufficient, meaning that smaller amounts of energy cannot cause considerable deformation of spacetime. This aligns with our understanding of gravity.
Acknowledgements: The authors are grateful to Professor J. W. Maluf for his invaluable assistance in revising the manuscript and for providing insightful ideas that greatly contributed to this work, and the author B.C.C. Carneiro is grateful to R. C. Puraca for assistance in Figure 1.